OpenStudy (jayden):

Write the equation of a line in slope intercept form that passes through the points (-2, 7) and (3, 12). Do not use spaces and enter fractions as improper reduced fractions when needed. is the right answer y=x-9

OpenStudy (anonymous):

no the sign is wrong. check your working

OpenStudy (jayden):

so it is y=x+9

OpenStudy (anonymous):

yes

OpenStudy (jayden):

\[3=12+b 3-12=12-12+b\] wouldn't 3-12 equal -9

OpenStudy (anonymous):

the answer is y = x + c subst in (-2,7) 7 = -2 + c c = 9

OpenStudy (anonymous):

thus y = x + 9

OpenStudy (anonymous):

if you use the other point, subst in (3,12) 12 = 3 + c c = 9 thus it is still y = x + 9

OpenStudy (anonymous):

1) Find your slope m= (y2-y1/x2-x1) = (12-7/3-(-2)) = 1 2) sub in one of your points to the equation y=mx+b, solve for b 12 = 1(3)+b 12-3=b 9= b 3) therefore, your equation is y=x+9

OpenStudy (jayden):

is that the slop intercept formula?

OpenStudy (jayden):

ow ok

OpenStudy (jayden):

Write an equation of a line in slope intercept form that is parallel to y = 3x+6 and passes through the point (-10, 2.5). so the answer for this q would be y=3x-32.5

OpenStudy (anonymous):

yep that's correct the answer is y=3x-32.5 :)

OpenStudy (jayden):

are you sure cause i have a vhs teacher saying that it is wrong

OpenStudy (anonymous):

its wrong

OpenStudy (anonymous):

the slope is the same since its parallel. thats 3 so the equation of that line is: y = 3x+c

OpenStudy (anonymous):

substitute in the point (-10,2.5) 2.5=3(-10) + c

OpenStudy (anonymous):

c = 32.5

OpenStudy (anonymous):

its y = 3x + 32.5

OpenStudy (anonymous):

irkiz is correct, sorry!

OpenStudy (jayden):

were did you get that formula cause it does not say that in my lesson do you use that formula for slope intercept form and point slope form?

OpenStudy (anonymous):

equation of a straight line in general is y = mx + c m is the slope/gradient. c is the y-intercept

OpenStudy (anonymous):

if you have the gradient and any point on that line, just substitute in the values and you get c

OpenStudy (jayden):

it has it as y=mx+b

OpenStudy (anonymous):

c or b doesnt matter.... its just an arbitary alphabet

OpenStudy (jayden):

Write an equation of a line in slope intercept form that is perpendicular to y = -4x -2 and passes through the point (-16, -11). Do not use spaces and enter fractions as improper reduced fractions when needed. the answer is y=4x+53 correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

eh no

OpenStudy (anonymous):

hold on let me check i said it without checking

OpenStudy (anonymous):

yup its right

OpenStudy (jayden):

thank you

OpenStudy (jayden):

y=-4/3+5 and the points that i am using are (0,-6) so you do y-o=3/4(x-0) and you get y=3/4x right

OpenStudy (jayden):

no that is wrong it is y-0=3/4(x--6) and you get y=3/4x+5.25